# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ 4.3.4. Batch Model Reuse Mechanisms
Table 6. Comparative analysis of FL studies categorized by our taxonomy for batch model reuse mechanisms. Studies applicable and conditional applicable to query-based FL are marked with different colors; Purple denotes operations completed on Server, or knowledge distilled from Global or Consensus Model; Blue denotes operations completed on Clients, or knowledge distilled from Local, Personlized, or Generative Models; Knowledge or Generated Content based on Public, Proxy or Generated data, and Knowledge or Generated Content based on Local, Private or Sensitive data; [ ]*1: one round of communication (aka one-shot), [ ]*N: multiple rounds of communication, processes ahead …[ ] are performed only once (i.e. preprocessing), processes inside […] are main functional part; Slash ”/”: model training based on private data, Comma ”,”: model training based on non-sensitive data; Goals of works: Efficiency Heterogeneity Privacy.
@@
Contrastive Learning Logits Avg Continual Learning Logits
@@
NeighGen (Zhang et al., 2021c)
FedSage (Zhang et al., 2021c)
@@
KD Hidden, Attention, Logits KD Hidden, Attention, Logits
@@
KD Softmax KD Softmax, Interpolated Softmax
@@
KD Logits Avg, Batch-Wise Statistics, Softmax
KD Softmax, Softmax
@@
KD Softmax, Softmax Avg KD Softmax Avg KD Softmax
@@
4.3.5. Hybrid Model Reusing in FL
Following the taxonomy we introduced in Section 4.3.4, it can be observed that almost all FL studies can be regarded as a permutation of four model reuse mechanisms: Combination, Amalgamation, Distillation, and Generation. To enhance our understanding of current model reusing studies and identify methods applicable for constructing an open FL platform, we provide a comprehensive summary in Table 7. We employ different colors and fonts in Table 7 to emphasize the distinctions among studies, while certain processes have been omitted without ambiguity. In addition, we have listed the main goals of each study, and only those with explicit designs, experiments, or proof are counted. Please refer to the table caption for the explanation of our denotations. As an example, the process of RSCFed (Liang et al., 2022) involves the following steps: ① KD, the knowledge is extracted as the softmax values from the self model based on local private data (ref. Table. 4) and performed at the client-side; ② Model exponential moving averaging performed at the client-side; ③ Simple Model Averaging performed two times at the server-side. These processes are repeated across multiple communication rounds until completion. In this way, we can categorize these works and make intuitive comparisons between them. We demonstrate the usage of our taxonomy by presenting three novel findings based on it.
Finding 1. Amalgamation and Distillation ([AD] or [DA]) are the most popular combo in FL studies, followed by Distillation as a standalone ([D]) strategy and Generation before Amalgamation ([GA]).
Due to the presence of parameter mismatch (Yu et al., 2021; Wang et al., 2020b; Yurochkin et al., 2019a), using solely amalgamation often lead to weight divergence (Li et al., 2020d; Tan et al., 2022b), particularly in scenarios with high data skew. As a result, distillation, which provides a global view of the data distribution, serves as a complementary solution for addressing data heterogeneity (indicated as ”H” in Table 7). However, to achieve acceptable performance, amalgamation requires multiple accesses to local data for training and multiple communications for model averaging ([AD]*N or [DA]*N), which makes it challenging to apply in one-shot FL settings.
On the other hand, we can fuse the distilled knowledge for training without any model amalgamation ([D]*N or [D]*1). However, additional precautions should be taken to prevent sensitive knowledge leakage. One common approach is to introduce an auxiliary public unlabeled dataset for knowledge extraction (except for FedED (Sui et al., 2020), which uses labeled validation set, ref. Table 4). Getting rid of the limitation of amalgamation, these distillation studies (e.g., FedAD (Gong et al., 2021), FedKD1 (Gong et al., 2022)) enable the feasibility of one-shot learning, making them good candidates for constructing a query-based FL platform. Meanwhile, the generation approachs also offer the same one-shot capacity (i.e., FedDISC (Yang et al., 2023) and FRD (Cha et al., 2019)) as distillation by directly generating synthetic data for fusion. This one-shot capacity can also be extended to the hybrid case, such as FedCAV (Heinbaugh et al., 2023), FedBE (Chen and Chao, 2021), FedAUX (Sattler et al., 2021), DENSE (Zhang et al., 2022a). That is why the final step in all one-shot solutions is typically distillation or generation on the server-side (excluding local training), rather than amalgamation. But it is worth noting that the one-shot feasibility is not sufficient and necessary in the context of query-based FL. For example, in FRD (Cha et al., 2019), participants generate mixup data for server training rather than contributing their models. Currently, it is evident that the once strong connection between model averaging and FL has become blurred, and distillation has emerged as a popular method in the field of FL.
Finding 2. Combination is the least common model reusing method in FL studies.
One drawback of using combination in the FL context is the high communication consumption required for broadcasting the combined model to multiple participants, which can be expensive in terms of bandwidth and latency. Hence, a practical approach is to compress the contributed models through methods like amalgamation or distillation, which may already implicitly include combination. For example, in FedAvg, the FL server collects local models from clients and averages them immediately. However, to avoid redundancy, we did not include these temporary combinations in Table 7.
Finding 3. Amalgamation after Distillation ([DA]) primarily distills knowledge from private data, while Distillation after Amalgamation ([AD]) and solely Distillation ([D]) primarily distill knowledge from non-sensitive data.
This difference arises from two different FL agreement mechanisms. One strategy is to leverage model amalgamation to achieve agreement ([DA]). For example, the knowledge from multiple data sources is generalized by applying appropriate weighting before being transferred to local models, which are then averaged to achieve the target model. Another strategy is to leverage a public dataset to achieve agreement ([D] or [AD]). In this approach, the knowledge distilled from multiple local models can be fused to a consensus agreement and then applied to the target model, eliminating the need for model amalgamation.
Question: How to select the suitable model reusing method for query-based FL?
From the previous discussion, we have demonstrated how to summarize recent FL studies from a model reusing perspective. Similarly, to select a suitable method for query-based FL, we can refer to Table 7 and consider the following factors:
  • Data Dependency. Recalling that query-based FL is contactless (Section 4.1), which means we cannot access the training data again once the model has been uploaded by the entity. Therefore, any method that requires multiple communication rounds of local or sensitive data access is not applicable to query-based FL. Such as, local data dependency due to local training (FedMD (Li and Wang, 2019), DynaFed (Pi et al., 2022), etc.), distillation (FedFusion (Yao et al., 2019), MOON (Li et al., 2021b), etc.), generation (SemiFL (Diao et al., 2022), NeighGen (Zhang et al., 2021c), etc.). So, refer to Table 7, the filtering criteria for unqualified methods can be expressed as […/…]*N, […D…]*N with knowledge in italic and […G…]*N with generated content in italic.
  • User Privacy. Due to the openness of the system, directly sharing embeddings from a pre-trained encoder or mixup of local data poses a risk of leakage (FedDISC (Yang et al., 2023), FRD (Cha et al., 2019)). It is strongly discouraged to include such methods, even if they have one-shot feasibility.
  • Compatibility. Additional, since query-based FL is model agnostic, it is not always possible to guarantee the compatibility of coordinate-wise operations required by model amalgamation. Therefore, the applicability of methods that involve model amalgamation may be limited in this context. We call these methods are conditional applicable to query-based FL.
In summary, we have marked the applicable and conditionally applicable methods to query-based FL in Table 7 in red and orange, respectively. Even though the KD procedure of DENSE (Zhang et al., 2022a) relies on batch-wise statistics collected during local training, we consider this information to be non-sensitive. Therefore, we include it in the applicable methods. Please note that some single combination or amalgamation approaches, such as Fed-ensemble (Shi et al., 2023) (only combination) and Models Fusion (Yurochkin et al., 2019b, a; Lam et al., 2021) (only amalgamation), can be directly applied in a query-based FL setting. However, to avoid redundancy, these studies have been omitted from Table 7. By now, we have witnessed the success of using the new taxonomy, and in the next section, we will discuss how to combine our taxonomy with the context of model licenses.
4.3.6. Model Reusing vs. Model Licenses
We have demonstrated that almost all FL studies can be regarded as a permutation of model reusing methods. Therefore, we can further analyze the corresponding clauses in model licenses perspective for FL studies by decomposing them into model reuse methods. Recalling that there are four kinds of model reusing mechanisms: Combination, Amalgamation, Distillation, and Generation, which result in four forms of outputs: Combination with strong separation, Combination with weak separation, Derivative from learned concepts, and Derivative from learned distributions, respectively. By interpreting ”work” as ”model” rather than as ”software”, we can easily identify the applicable clauses for different model reusing mechanisms.
We present the analysis of model licenses for model reuse in Table 8. The analysis covers the most popular model licenses (used by over 100 released models) on Hugging Face, which include free software licenses, free content licenses, and AI model licenses. For simplicity, we have merged some similar licenses without loss of clarity. To avoid license conflicts when engaging in batch model reuse, it is preferable for the reused model to be considered an independent work or fall under an undefined category, not covered by the original license. Additionally, we can refer to Table 8 and consider the following factors:
  • Differences in terminology. In the definitions of licenses, connecting multiple works into a separable union is typically referred to as aggregation, redistribution, and reproduce, which is completely different from the concept of model aggregation in the ML field. Therefore, it is necessary to refer to Table 8 for the corresponding delineations of each model reuse mechanism instead of relying solely on the technical name.
  • Restrictions of combination and amalgamation. As shown in Table 8, the majority of model licenses have specific terms and definitions regarding these two methods, except for blackbox combination. This implies that there may be potential restrictions that need to be taken into account and complied with. These restrictions may also proliferate to the reused results if original licenses of models are copyleft, such as GPL-3.0, CC-BY-SA, CC-BY-NC-SA, etc. Therefore, we recommend avoiding the use of models and datasets under such copyleft licenses (ref. Fig. 4(b) and Tabel 3) for batch model combination and amalgamation.
  • User behavioral restrictions of responsible AI licenses. OpenRAIL licenses (Contractor et al., 2022) clearly define all four model reuse mechanisms, which means the effect of copyleft-style behavioral-use clauses will spread to the reused results, which can potentially result in the licensed artifacts becoming closed source (Greenbaum, 2016) (ref. Section 4.3.2). To prevent license proliferation (Gomulkiewicz, 2009), it is not recommended to reuse models under these OpenRAIL licenses.
Table 8. Analysis of the model licensing clauses corresponding to different batch model reuse mechanisms, denoted by ”keywords in licenses” -¿ ”delineation of reused results”, ✗: undefined. Note: GPL-3.0, CC-BY-SA, $C$-BY-NC-SA are copyleft.
Combination Amalgamation Distillation Generation
Combinated Work
with Strong Separation
Combinated Work
with Weak Separation
Derivative Work
from Concepts
Derivative Work
from Distributions
Apache-2.0
Separable -¿ Independent Work
Modify -¿ Derivative
MIT
AFL-3.0
Modify -¿ Derivative
GPL-3.0
Blackbox: Aggregate -¿
Independent Work
Other: Link -¿ Modified Version
Modify -¿ Covered Work
Output no constitutes a covered work -¿
Independent Work
Output no constitutes a covered work -¿
Independent Work
Artistic-2.0
Blackbox: Merely Extend -¿
Own Work
Other: Link -¿ Own Work
Aggregate -¿
Modified Version
BSD-3-Clause
Blackbox: Rredsitribution in
binary forms -¿ ✗
Other: Redistribution of source code -¿ ✗
WTFPL-2.0
OpenRAIL Licenses
Transfer of patterns of
output -¿ Derivative
Transfer of patterns of weights -¿ Derivative
Transfer of patterns of activations -¿ Derivative
Transfer of patterns of output -¿ Derivative
Creative Commons
Licenses
Reproduce -¿ Adapted Material
Adapt -¿
Adapated Material
CC0-1.0
Reproduce -¿
Independent Work
Adapt -¿
Independent Work
Lastly, we can further analyze the potential license conflicts in hybrid model reuse methods. Let’s consider the example of DENSE (Zhang et al., 2022a), where the process is /C[G,DGD]*1 and assume that the local models are licensed under GPL-3.0. In this case, the collection of local models should be treated as a whitebox in order to distill the logits value from it. As a result, this collection should be considered a modified version of the original works according to the terms of GPL-3.0. However, the generation and distillation processes of DENSE are based on non-sensitive data and result in the creation of an independent work. Considering that GPL-3.0 is a copyleft license, the reused results should be licensed under GPL-3.0 if the new release includes the collection in its whitebox form. Otherwise, if the GPLed local models executed as a separable and replaceable subprogram, the reused results can be considered an independent work without any specific license restrictions.
The example above demonstrates that analyzing hybrid methods under a copyleft license can be complicated. Therefore, selecting models, algorithms, datasets with more permissive licenses, as advocated in Section 4.3.2, can make things easier and facilitate the model reusability of open FL platforms.
4.4. How to Protect Models
Figure 5. Overview of DNNs IP protection methods.
[…]
In the previous sections, we have extensively discussed the construction of a query-based FL platform from both technical and legal perspectives. Nonetheless, the matter of model management and protection continues to be a significant concern, raising questions such as How can I protect my models from plagiarism after they are released? In general, there are three kinds of mechanisms that can be employed to protect DNN models: authorization (Hanzlik et al., 2021), watermarking (Tekgul et al., 2021), and fingerprinting (Li et al., 2023b). Authorization mechanisms, including hardware-based memory encryption techniques like TEE, can effectively prevent unauthorized access to the model copy. Additionally, researchers have explored embedding authorization mechanisms within the DNN architecture itself, such as with passport layers (Fan et al., 2019; Zhang et al., 2020). However, the need for specific hardware and modifications to model architectures can significantly increase the barriers for participants and compromise the openness of query-based FL platforms.
As passive protection methods, watermarking and fingerprinting can detect potential infringement while preserving the openness of the model. A watermark can be embedded into DNNs through training or fine-tuning with a modified loss function (Uchida et al., 2017) or trigger data (Darvish Rouhani et al., 2019) acting as a backdoor (Bagdasaryan et al., 2020). On the other hand, fingerprinting strategies aim to generate an identifier for each model based on snapshots of training history (Jia et al., 2021) or inference results that approximate decision boundaries (Chen et al., 2022). Note that the technical details of IP protection methods are beyond the scope of this survey. We recommend referring to up-to-date surveys such as (Peng et al., 2022; Sun et al., 2023). Additionally, it is important to recognize that the goal of query-based FL platforms is not to provide perfect protection against model plagiarism and unauthorized use. Interestingly, certain model reusing strategies like KD and ensemble methods are even regarded as model extraction attacks within the realm of DNNs IP protection (Charette et al., 2022). This highlights the inherent contradiction between promoting model reusing and preventing all forms of model plagiarism. As a result, it is necessary to determine the scope of protection and strike a balance between model protection and platform openness.
First, it is helpful to consider the following requirements for model protection in the context of query-based FL:
  • Non-invasive. Any attempt to invasively embed backdoors into the weights and architectures of models is likely to result in changes to their functionalities and could lead to unexpected failures after deployment (Chen et al., 2022). Meanwhile, these backdoors can be exploited by attackers to manipulate the model’s predictions (Li et al., 2022a). Accordingly, it is more recommend leaving the decision of adding invasive protection to the users instead of forcing it through platforms.
  • Compatible. Due to the model agnostic nature of query-based FL, relying on protection mechanisms that are specific to certain model structures and formats will limit the applicability of the platform. For instance, the passport layer (Zhang et al., 2020) is built upon the normalization layer and requires joint training with the target model for ownership verification. However, in practice, both the normalization layer and whitebox access to the model may be unavailable. Therefore, considering the compatibility of model protection mechanisms is essential to ensure wide support for model sharing in query-based FL.
  • Permissive. As mentioned earlier, it is not essential to identify and address every instance of plagiarism in query-based FL. On the contrary, we encourage platform users to engage in model mining and reusing, as outlined in Section 4.3.4. Therefore, model protection methods should be permissive enough to allow for model reusing while still identifying the less creative effort operations that result in minimal or no change to the model’s functionality, such as naive replication, quantization, pruning, and invariant neural swapping. Note that the presence of sufficient human creative effort is an important criterion for determining the copyrightability of a computer-generated work (on New Technological Uses of Copyrighted Works, US). The consideration of copyrightability can serve as a guiding principle in determining the scope of model protection in query-based FL.
  • Large-scale cost-effective. Unlike traditional FL, where the size of training networks in each round is fixed, the size of a query-based FL platform is continuously expandable. Therefore, it is undesirable if the potential conflicts (Li et al., 2023b), required bit-length, and deployment cost (Uchida et al., 2017; Darvish Rouhani et al., 2019) of model protection solutions such as watermarking and fingerprinting increase with the platform size. This suggests that solutions that require modifying optimization goals and training or fine-tuning with trigger data to embed identity information into the target model should be excluded.
Recently, several model protection algorithms have been proposed to address the challenge of model protection in FL. Tekgul et al. (Tekgul et al., 2021) propose WAFFLE that uses augmented Gaussian noisy images as a watermark set to determine the ownership of the global model. However, this method is invasive and incurs high costs due to multiple rounds of training. To ensure reliable watermarking for FL participants, FedIPR (Li et al., 2023b) embeds different watermarks into models of different participants through training with additional trigger samples and a modified optimization goal. But this method is designed for model verification on the client-side, so it is independent to the model management in FL platforms. FedTracker (Shao et al., 2022) embeds personalized local fingerprints into the Batch Normalization (BN) layer to enable traceability of model leakers. One limitation of FedTracker is that its fingerprinting strategies are not compatible with other DNNs that do not use BN or do not provide whitebox access. These examples highlight the incompatibility between the IP protection methods designed for traditional FL and the requirements of the query-based FL scenario, which have different protection goals and targets.
Based on the above observation and the taxonomy of deep IP protection (Peng et al., 2022; Sun et al., 2023), dynamic fingerprinting strategies with blackbox verification support are considered suitable model protection methods for query-based FL. These methods aim to approximate the similarities between the decision boundaries of different models by evaluating the value of a well-tailored testing metric on a self-constructed testing set. A typical example is DeepJudge (Chen et al., 2022), which employs an ensemble of multi-level (Property, Neuron, Layer) testing metrics to achieve confident and robust plagiarism identification. The verification process of DeepJudge can be evaluated in a blackbox setting by calculating distances based on model predictions. However, the generation of testing set in DeepJudge requires whitebox access to the suspect models in order to calculate adversarial samples, which may be unreachable if the suspect models are in binary format.
Another example is Zest (Jia et al., 2022), which randomly samples several training images and then applies super-pixel and segmentation techniques to construct the testing set. However, this testing samples generation method is limited to images and may expose private training data. A compromise approach is to seek non-sensitive or desensitized testing set from users when they upload their models. Inspired by DeepJudge, we can develop an ensemble of multi-level testing strategies for query-based FL. In the first step, we can compare the hash codes (e.g. MD5, SHA-256) of the raw model weights or its binary execution to rapidly filter out cases of naive replication plagiarism. Then, in the second step, we can filter out cases of quantization or invariant neural swapping plagiarism by comparing the distance in predictions obtained from an out-of-distribution dataset. In addition, we can further evaluate the similarities between models using testing-based approaches such as DeepJudge and Zest, which can detect cases of pruning and direct KD plagiarism. Furthermore, if the workflow information is supports it, we can swiftly eliminate batch model reusing from being suspected of plagiarism. To enhance understanding, we present an overview of the typical methods used for protecting IP of DNNs in Fig. 5.
In summary, finding a balance between model protection and platform openness is indeed a challenge. It is crucial to carefully consider the trade-offs and explore alternative approaches that can provide a reasonable level of model protection without excessively compromising the openness and usability of query-based FL platforms.
4.5. Limitations of Query-based FL
In query-based FL systems, the processes of model production and reuse are decoupled to maximize the autonomy of each participant. However, this losse cooperation paradigm is no longer compatiable with online collaboration ML frameworks, which means that it cannot fully harness participants’ communication and computational resource to enhace the training performance. Therefore, even though the server-dominated cooperation frameworks like FedAvg have limitations, as described in Section 3.2, it is still worthwhile to provide support for its underlying distributed ML training methods to enhance the flexibility and compatibility of our open FL platforms. In the next section, we illustrate another cooperation framework named contract-based FL, which follows a mutual choice design philosophy similar to crowdsourcing platforms (Vaughan, 2018). It can serve as an extension of traditional FL and query-based FL.
5. Contract-based Federated Learning
Let us consider another open FL platform based on a contract-based cooperation framework, where cooperation can be built by publishing ML tasks and accepting ML collaboration requests. An overview of this platform is presented in Fig. 6, the major difference between traditional FL (ref Fig. 2) and contract-based FL is that we involve a trustworthy third-party platform called the Contract Platform to host and coordinate ML tasks for platform users. Rather than directly pushing ML tasks from servers to clients, employers need to publish the tasks to the contract platform and wait for acceptance from workers, which involves a mutual choice procedure. The conditions and payment for participation are also appliable, and the final model will be audited and evaluated by the platform for fairness. In addition, privacy-enhanced technologies (Hesamifard et al., 2018) and fingerprint management (Chen et al., 2022) can also be implemented by the platform to protect the IP of platform users. The properties of contract-based FL are: 1) Opt-in, as workers reserve the right to join or quit from training networks; 2) Contractual, enabling employers to define payment, organization mode, model quality criteria, rehire rules, etc. through contracts; 3) Market-based, where contracts are open and task pricing is influenced and determined by the market.
In short, contract-based FL retains all functionalities related to model training at the platform and only provides task publishing services to employers, who play the role of model training in traditional FL. Furthermore, similar to model-reusing mechanisms (ref. Section 4.3.4) that can be applied to query-based FL, our contract-based FL is not limited to an amalgamation-based collaborative training paradigm like FedAvg. In fact, the contract platform is designed as a crowdsourcing platform among data miners for ML cooperation tasks, including distributed training, fine-tuning, and ensembling based on scratch or Foundation Models (FMs) (Yuan et al., 2022). However, most current crowdsourcing platforms like Amazon Mechanical Turk [^20] https://www.mturk.com/ and Appen [^21] https://appen.com/crowd-2/ primarily deal with human intelligence tasks such as data collection and annotation (aka. microtasks) rather than ML modeling tasks. It remains unexplored how to design and publish an ML task to seek crowd labor. On the other hand, many studies (Dias and Meratnia, 2022; Blythman et al., 2022; Deng et al., 2021; Guo et al., 2023; Batool et al., 2022) and products (Steeves et al., 2022; Ziller et al., 2021; McConaghy, 2022) emphasize monetizing ML activities through blockchain-based techniques and Decentralized Autonomous Organizations (DAOs). However, both monetization and decentralization are non-essential functions for our contract platforms, and we will summarize and compare these works later.
Figure 6. An overview of contract-based FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
[…]
Table 9. Comparisons of typical model training frameworks with the intervention of a trustworthy platform. *, *, * indicate resources from Employer, Platform, and Worker, respectively. Resources in different color columns (e.g., Data in column Platform) represent transmission. $\bigstar$: Location of model training. Resources received from other workers.

### 29 Feb 2024 v3

@@ V-E Hybrid Model Reusing and Model Licenses
TABLE V: Comparative analysis of FL studies categorized by our taxonomy for batch model reuse mechanisms. Studies applicable and conditional applicable to query-based FL are marked with different colors; Purple denotes operations completed on Server, or knowledge distilled from Global or Consensus Model; Blue denotes operations completed on Clients, or knowledge distilled from Local, Personlized, or Generative Models; Knowledge or Generated Content based on Public, Proxy or Generated data, and Knowledge or Generated Content based on Local, Private or Sensitive data; [ ]*1: one round of communication (aka one-shot), [ ]*N: multiple rounds of communication, processes ahead …[ ] are performed only once (i.e. preprocessing), processes inside […] are main functional part; Slash "/": model training based on private data, Comma ",": model training based on non-sensitive data; Goals of works: Efficiency Heterogeneity Privacy.
@@
Contrastive Learning Logits Avg
Continual Learning Logits
@@
NeighGen [197]
FedSage [197]
@@
KD Hidden, Attention, Logits
KD Hidden, Attention, Logits
@@
KD Softmax
KD Softmax, Interpolated Softmax
@@
KD Logits Avg, Batch-Wise Statistics, Softmax KD Softmax, Softmax
@@
KD Softmax, Softmax Avg
KD Softmax Avg KD Softmax
@@
VI Remaining Topics in Query-based FL
VI-A Model Protection
In the previous sections, we have extensively discussed the construction of a query-based FL platform from both technical and legal perspectives. However, model management and protection continues to be a significant concern, raising questions such as How can I protect my models from plagiarism after they are released? Our conclusion is using dynamic fingerprinting strategies with blackbox verification support‡V, such as DeepJudge [211] and Zest [212].
VI-B Limitations of Query-based FL
In query-based FL systems, the processes of model production and reuse are decoupled to maximize the autonomy of each participant. However, this losse cooperation paradigm is no longer compatiable with online collaboration ML frameworks, which means that it cannot fully harness participants’ communicational and computational resource to enhace the training performance. Therefore, even though the server-dominated cooperation frameworks like FedAvg have limitations, as described in §III-B, it is still worthwhile to provide support for its underlying distributed ML training methods to enhance the flexibility and compatibility of our open FL platforms. In the next section, we illustrate another cooperation framework named contract-based FL, which follows a mutual choice design philosophy similar to crowdsourcing platforms [213]. It can serve as an extension of traditional FL and query-based FL.
VII Contract-based Federated Learning
VII-A Overiew
Let us consider another open FL platform based on a contract-based cooperation framework, where cooperation can be built by publishing ML tasks and accepting ML collaboration requests. An overview of this platform is presented in Fig. 6, the major difference between traditional FL (ref Fig. 3) and contract-based FL is that we involve a trustworthy third-party platform called Contract Platform to host and coordinate ML tasks for platform users. Rather than directly pushing ML tasks from servers to clients, employers need to publish the tasks to the contract platform and wait for acceptance from workers, which involves a mutual choice procedure. The conditions and payment for participation are also appliable, and the final model will be audited and evaluated by the platform for fairness. In addition, privacy-enhanced technologies [42] and fingerprint management [211] can also be implemented by the platform to protect the IP of platform users. The properties of contract-based FL are: 1) Opt-in, as workers reserve the right to join or quit from training networks; 2) Contractual, enabling employers to define payment, organization mode, model quality criteria, rehire rules, etc. through contracts; 3) Market-based, where contracts are open and task pricing is influenced and determined by the market.
Figure 6: An overview of contract-based FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
In short, contract-based FL retains all functionalities related to model training at the platform and only provides task publishing services to employers, who play the role of model training in traditional FL. Furthermore, similar to model-reusing mechanisms (ref. §V) that can be applied to query-based FL, our contract-based FL is not limited to an amalgamation-based collaborative training paradigm like FedAvg. In fact, the contract platform is designed as a crowdsourcing platform among data miners for ML cooperation tasks, including distributed training, fine-tuning, and ensembling based on scratch or Foundation Models (FMs) [214]. However, most current crowdsourcing platforms like Amazon Mechanical Turk and Appen [^4] AMT: https://www.mturk.com/; Appen: https://appen.com/crowd-2/ primarily deal with human intelligence tasks such as data collection and annotation (aka. microtasks) rather than ML modeling tasks. It remains unexplored how to design and publish an ML task to seek crowd labor. On the other hand, many studies [215, 216, 217, 218, 219] and products [220, 58, 221] emphasize monetizing ML activities through blockchain-based techniques and Decentralized Autonomous Organizations (DAOs). However, both monetization and decentralization are non-essential functions for our contract platforms‡VI.
TABLE VI: Comparisons of typical model training frameworks with the intervention of a trustworthy platform. *, *, * indicate resources from Employer, Platform, and Worker, respectively. Resources in different color columns (e.g., Data in column Platform) represent transmission. $\bigstar$: Location of model training. Resources received from other workers.